A jet engine releases 5000j of energy per cycle and performs 800j work.


1. How much heat is absorbed by this engine per cycle.


2. What is the thermal efficiency


3. How much work this engine perform in a cycle


4. If this engine complete each cycle in 20sec, what is the power rating of this engine.

1
Expert's answer
2021-11-15T10:51:41-0500

1. 1-(800÷5000)=0.84j

2. 800÷5000=0.16j

3. 5000-800= 4200j

4. 4200÷20=210w